# Enhancing Fault Detection for Large Language Models via Mutation-Based Confidence Smoothing

### v1

@@ 3.5. Configurations
Configuration of fault detection methods. NNS and TestRank need to compute the distance between data samples and require the representation of the samples. In this work, we directly use the input embeddings extracted from LLMs as the representation. Besides, TestRank is a learning-based method where training data is required for fault detection. For the clone detection, problem classification, and TagMyNews tasks, we randomly select 200 data from the original datasets as the training data. For the sentiment analysis dataset, as the original dataset only contains 200 data samples and 150 of them are used as test data, we collect the remaining 50 samples as the training data. For the data labeling process, we set the labeling budgets as 10%, 30%, and 50%. Configuration of LLMs. In our experiments, we set top_p=0.9 for LLaMA/CodeLLaMA and top_p=1 for GPT3.5/GPT4 which are the default settings used by LLMs. top_p controls the accumulation of token probabilities to select the next word for all LLMs. We set the window size of the token context as max_length= 10k for LLaMA/CodeLLaMA to match our test data’s token length. The window sizes of GPT3.5 and GPT4 are set as 16385 and 128000 respectively, using the default settings.
@@ 3.6. Implementation and Environment
For the fault detection methods, we reuse the official implementations provided by the original papers and modify them to fit our tasks. For the LLaMA and CodeLLaMA models, we use the resources provided by Hugging Face 333Hugging Face, https://huggingface.co/ in this work. For GPT3.5 and GPT4 models, we use the OpenAI’s API to access the models and collect the results. In total, it costs 267 US dollars to run GPT-related experiments. We conduct all LLaMA-related experiments on a 2.6 GHz Intel Xeon Gold 6132 CPU with an NVIDIA Tesla V100 16G SXM2 GPU.
@@ 4.1. RQ1: Prediction Confidence
First, we check whether LLMs can confidently predict the data in our studied tasks. Figure 1 depicts the distribution (number of data in each confidence interval) of the prediction confidence produced by LLMs, where we set the number of intervals as 30. The first conclusion we can draw is that the series of GPT models are more confident with their predictions than the series of LLaMA models. On average, the confidence of LLaMA, CodeLLaMA, GPT3.5, and GPT4 on these four datasets are 0.5338, 0.6993, 0.7701, and 0.8727, respectively. Comparing the accuracy of LLMs and their average confidence (which is a notion of miscalibration (Guo et al., 2017)), we found that existing LLMs are not well-calibrated. For example, for CodeLLaMA on the clone detection task, the accuracy and average confidence are 49% and 0.758, respectively. This means CodeLLaMA is overconfident with its prediction even though it has poor performance. Overall, except for sentiment analysis, LLMs are overconfident in the other three tasks with 21.36% higher confidence than the true accuracy. Besides, ECE demonstrated that GPT4 is better calibrated than other LLMs in all our considered tasks. The Reliability diagrams (Guo et al., 2017) that visualize the model calibration can be found in Appendix A. Additionally, the results show that prediction confidence is concentrated in certain ranges for many cases. For example, for the clone detection - GPT3.5, confidence only falls into four ranges. This means the model has similar confidence to many different inputs, in turn, the output-based fault detection methods (e.g., MaxP) could produce similar uncertainty scores to these inputs.
Considering the other two cases that have poor performance reported in Table 1, CodeLLaMA for problem classification and LLaMA for TagMyNews. We found they have confused confidence on these tasks as shown in Figure 5(b) and Figure 5(c). For instance, in Figure 5(b), the most frequent confidence score is 0.5, and in Figure 5(c), there are no confidence scores that are higher than 0.8. This phenomenon indicates that CodeLLaMA and LLaMA cannot handle problem classification and news classification tasks at all.
Finally, we found that GPT4 has perfect accuracy (100%) on the problem classification task with high average confidence (0.9827 on average) and a low ECE score of 0.0143. We conjecture there is a data leakage problem of our test data, i.e., our used problem classification dataset is a part of the training data of GPT4. As there are no faults in GPT4 for the problem classification task, we exclude this model from the following fault detection study.
@@ 4.2. RQ2: Effectiveness of Fault Detection
In this RQ, we explore the performance of existing fault detection methods on LLMs. As mentioned in Section 3.3, we exclude models that have poor performance (nearly random guesses) and perfect performance (100% accuracy) on the studied tasks.
Budget | Random | Gini | Entropy | MCP | MaxP | Margin | ATS | NNS | TestRank | |
Sentiment Analysis | 10% | 0.0667 | 0.0667 | 0.0667 | 0.1333 | 0.0000 | 0.1333 | 0.0667 | 0.0000 | 0.0667 |
30% | 0.4375 | 0.0625 | 0.1250 | 0.5000 | 0.0000 | 0.1250 | 0.1250 | 0.1250 | 0.0625 | |
50% | 0.4375 | 0.3125 | 0.2500 | 0.5625 | 0.2857 | 0.4375 | 0.3750 | 0.1875 | 0.2500 | |
Average | 0.3139 | 0.1472 | 0.1472 | 0.3986 | 0.0952 | 0.2319 | 0.1889 | 0.1042 | 0.1264 |
Budget | Random | Gini | Entropy | MCP | MaxP | Margin | ATS | ||
GPT3.5 | Clone Detection | 10% | 0.1500 | 0.3000 | 0.4500 | 0.3500 | 0.3000 | 0.3000 | - |
30% | 0.3265 | 0.4694 | 0.4490 | 0.4082 | 0.4694 | 0.4694 | - | ||
50% | 0.5102 | 0.6327 | 0.6122 | 0.6122 | 0.6327 | 0.6327 | - | ||
Problem Classification | 10% | 0.7000 | 0.8000 | 0.7500 | 0.7000 | 0.8000 | 0.8500 | 0.6500 | |
30% | 0.7167 | 0.7667 | 0.6500 | 0.6333 | 0.7833 | 0.7667 | 0.6833 | ||
50% | 0.5900 | 0.7000 | 0.6100 | 0.6400 | 0.7000 | 0.7000 | 0.6700 | ||
TagMyNew | 10% | 0.7000 | 0.5500 | 0.5500 | 0.6000 | 0.5500 | 0.5500 | 0.4000 | |
30% | 0.4000 | 0.5333 | 0.5333 | 0.4500 | 0.5500 | 0.5500 | 0.4889 | ||
50% | 0.4375 | 0.5625 | 0.5500 | 0.5000 | 0.6000 | 0.6625 | 0.8444 | ||
Sentiment Analysis | 10% | 0.0667 | 0.2667 | 0.1333 | 0.2667 | 0.2667 | 0.3333 | 0.0000 | |
30% | 0.3704 | 0.3704 | 0.4074 | 0.4444 | 0.4444 | 0.4074 | 0.2222 | ||
50% | 0.5926 | 0.6667 | 0.7037 | 0.5556 | 0.6667 | 0.7037 | 0.5185 | ||
Average | 0.4634 | 0.5515 | 0.5333 | 0.5134 | 0.5636 | 0.5771 | 0.4975 | ||
GPT4 | Clone Detection | 10% | 0.4000 | 0.5000 | 0.5000 | 0.2500 | 0.5000 | 0.2000 | - |
30% | 0.3478 | 0.5652 | 0.5435 | 0.5217 | 0.5652 | 0.3478 | - | ||
50% | 0.5435 | 0.6522 | 0.6522 | 0.6739 | 0.6522 | 0.5652 | - | ||
TagMyNews | 10% | 0.3000 | 0.4500 | 0.3500 | 0.4000 | 0.4500 | 0.4500 | 0.4000 | |
30% | 0.3333 | 0.6667 | 0.6667 | 0.5000 | 0.6389 | 0.6667 | 0.7222 | ||
50% | 0.5278 | 0.8611 | 0.8611 | 0.7778 | 0.8611 | 0.8611 | 0.8333 | ||
Sentiment Analysis | 10% | 0.0667 | 0.3333 | 0.5333 | 0.3333 | 0.3333 | 0.3333 | 0.2667 | |
30% | 0.4667 | 0.6667 | 0.7333 | 0.7333 | 0.6667 | 0.6667 | 0.3333 | ||
50% | 0.6000 | 0.9333 | 1.0000 | 0.9333 | 0.9333 | 0.9333 | 0.5333 | ||
Average | 0.3984 | 0.6254 | 0.6489 | 0.5693 | 0.6223 | 0.5582 | 0.5148 |
Table 2 and Table 5 summarize the results. The first conclusion we can draw from the results is that the existing fault detection methods cannot significantly find faults in LLMs compared to their performance on classical deep learning models (LI et al., 2021) where the best method can achieve a 0.9532 TRC score. For example, considering the detection performance on the LLaMA model, only the best method has a higher TRC score (0.3986) than random selection (0.3139). This means all the other well-designed methods perform worse than random guesses. Considering the results on GPT3.5 and GPT4 models, this phenomenon has been alleviated. All the methods perform better than random selection. However, the best methods on GPT3.5 and GPT4 have only 0.5771 and 0.6489 TRC scores. There is still a big gap between the existing performance and the idea performance, TRC=1.
Then, comparing each method, we found that no method can consistently perform better than the others. The performance of fault detection methods on LLMs highly depends on the datasets and models. However, it is interesting that when the performance of LLMs becomes better, the fault detection methods become more effective. For example, the TRC scores produced by each method on GPT4 (on average 0.6114) are much higher than the TRC scores on LLaMA (on average 0.3074). The reason is that existing methods that are based on output probabilities highly rely on the confidence of models in their outputs. As GPT4 is more confident that LLaMA, the fault detection methods can perform better on GPT4 accordingly.
Answer to RQ2: Existing fault detection methods are not effective in detecting faults in LLMs and there are no methods that can consistently perform better than the others. Methods have better performance on LLMs that have higher accuracy (GPT4) than those that have lower accuracy (GPT3.5 and LLaMA).
5. MuCS
@@
Since the concentrated confidence will harm the performance of output-based fault detection methods (those methods will give similar uncertainty scores to the inputs), we need to diversify the prediction confidence of LLMs to enhance fault detection. It is challenging since we cannot get the internal information (e.g., output logits) of closed-source LLMs to smooth the prediction confidence. Inspired by previous works (Wang et al., 2021; Li et al., 2023b) that showed the potential of testing deep learning models by mutation analysis (one uses mutation killing score to detect faults and the other one uses mutation analysis to improve the performance of code models), we propose MuCS, a simple yet effective black-box solution to enhance fault detection by smoothing the prediction confidence from the input-level using prompt mutation.
@@
Figure 2 presents the overall workflow of our proposed solution. It has two main steps, 1) prompt mutation, and 2) confidence smooth. Specifically, given an LLM and an input prompt p, we first utilize transformation methods to mutate the inputs and generate n mutated prompts, , , …, . For NLP tasks, we use text augmentation methods (Marivate and Sefara, 2020) (e.g., token random deletion) to transform prompts into new ones. For the code tasks, we consider both text augmentation methods and code refactoring methods (Al Dallal and Abdin, 2017) (e.g., add new Print() functions) for the prompt mutation. After that, we feed all the mutants to the LLM and collect the output confidence , ,…, produced by the LLM. Finally, a confidence smooth method is employed to compute the final confidence of the LLM on the input. In this work, we use the straightforward method, average, to smooth the confidence. Algorithm 1 summarizes our mutation-based confidence smoothing solution.
@@ 5.1. Mutation Operators
Two types of mutation operators have been used, text augmentation methods and code refactoring methods. In total, five text augmentation methods (including Synonym Replacement, Random Deletion, Random Insertion, Random Swap, and Punctuation Insertion), and four code refactoring methods (including Print Adding, Local Variable Adding, Dead If Adding, and Duplication) have been used in MuCS. Given a prompt with the length of works , each mutation operator is defined as follows.
@@ 6.1. Research Question
To check whether MuCS can help fault detection or not, we first analyze the effectiveness of existing fault detection methods using MuCS. Then, we explore the reason why MuCS enhances fault detection by analyzing its influences on the prediction confidence of LLMs.
RQ3: How effective is MuCS in enhancing fault detection? In this research question, we compare the effectiveness of fault detection methods with and without using MuCS to analyze the usefulness of our proposed solution.
RQ4: How does MuCS affect the outputs of the LLMs? We want to explore how MuCS influences the outputs of LLMs to check why it works on fault detection.
@@ 6.3. RQ3: Effectiveness of MuCS
Budget | Gini-M | Entropy-M | MCP-M | MaxP-M | Margin-M | ATS-M | NNS-M | TestRank-M | BALD |
10% | 0.0000( 100.00%) | 0.0667( 0.00%) | 0.1333( 0.00%) | 0.0000( * ) | 0.2667( 100.08%) | 0.0667( 0.00%) | 0.0000( * ) | 0.0000( 100.00%) | 0.2667 |
30% | 0.1250( 100.00%) | 0.1250( 0.00%) | 0.5000( 0.00%) | 0.2500( * ) | 0.3125( 150.00%) | 0.2500( 100.00%) | 0.0625( 50.00%) | 0.1250( 100.00%) | 0.5625 |
50% | 0.5625( 80.00%) | 0.5625( 125.00%) | 0.7500( 33.33%) | 0.5625( 96.88%) | 0.6250( 42.86%) | 0.3750( 0.00%) | 0.3750( 100.00%) | 0.5625( 125.00%) | 0.8125 |
Average | 0.2292( 26.67%) | 0.2514( 41.67%) | 0.4611( 11.11%) | 0.2708( 96.88%) | 0.4014( 97.64%) | 0.2306( 33.33%) | 0.1458( 25.00%) | 0.2292( 41.67%) | 0.5472 |
Budget | Gini-M | Entropy-M | MCP-M | MaxP-M | Margin-M | ATS-M | BALD | ||
GPT3.5 | Clone Detection | 10% | 0.4500( 50.00%) | 0.4500( 0.00%) | 0.4500( 28.57%) | 0.4500( 50.00%) | 0.4500( 50.00%) | - | 0.4500 |
30% | 0.5306( 13.04%) | 0.5102( 13.63%) | 0.5306( 29.99%) | 0.5306( 13.04%) | 0.5306( 13.04%) | - | 0.3265 | ||
50% | 0.6939( 9.67%) | 0.5510( 10.00%) | 0.5918( 3.33%) | 0.6939( 9.67%) | 0.6939( 9.67%) | - | 0.5714 | ||
Problem Classification | 10% | 0.8000( 0.00%) | 0.8000( 6.67%) | 0.5500( 21.43%) | 0.8500( 6.25%) | 0.8000( 5.88%) | 0.8500( 30.77%) | 0.8500 | |
30% | 0.8333( 8.69%) | 0.7833( 20.51%) | 0.6500( 2.64%) | 0.8500( 8.52%) | 0.8500( 10.86%) | 0.7667( 12.21%) | 0.8500 | ||
50% | 0.8000( 14.29%) | 0.8300( 36.07%) | 0.6900( 7.81%) | 0.8000( 14.29%) | 0.7700( 10.00%) | 0.7200( 7.46%) | 0.7700 | ||
TagMyNew | 10% | 0.6000( 9.09%) | 0.6500( 18.18%) | 0.4000( 33.33%) | 0.7000( 27.27%) | 0.6000( 9.09%) | 0.7000( 75.00%) | 0.7500 | |
30% | 0.6000( 12.51%) | 0.6000( 12.51%) | 0.4500( 0.00%) | 0.6167( 12.13%) | 0.5833( 6.05%) | 0.6042( 23.58%) | 0.6167 | ||
50% | 0.6875( 22.22%) | 0.6625( 20.45%) | 0.5750( 15.00%) | 0.7000( 16.67%) | 0.6750( 1.89%) | 0.7917( 6.24%) | 0.6500 | ||
Sentiment Analysis | 10% | 0.3333( 24.97%) | 0.4000( 200.08%) | 0.5333( 99.96%) | 0.2667( 0.00%) | 0.3333( 0.00%) | 0.4000( * ) | 0.3333 | |
30% | 0.5926( 59.99%) | 0.5926( 45.46%) | 0.7037( 58.35%) | 0.5556( 25.02%) | 0.6667( 63.65%) | 0.5926( 166.70%) | 0.5556 | ||
50% | 0.9259( 38.88%) | 0.8889( 26.32%) | 0.8148( 46.65%) | 0.9259( 38.88%) | 0.9259( 31.58%) | 0.6667( 28.58%) | 0.6296 | ||
Average | 0.6539( 21.95%) | 0.6432( 32.49%) | 0.5783( 19.24%) | 0.6616( 18.48%) | 0.6566( 16.66%) | 0.6769( 42.26%) | 0.6128 | ||
GPT4 | Clone Detection | 10% | 0.4500( 10.00%) | 0.4500( 10.00%) | 0.4000( 60.00%) | 0.4500( 10.00%) | 0.3000( 50.00%) | - | 0.2500 |
30% | 0.6304( 11.54%) | 0.6304( 15.99%) | 0.6304( 20.84%) | 0.6304( 11.54%) | 0.2391( 31.25%) | - | 0.2222 | ||
50% | 0.7826( 19.99%) | 0.7391( 13.32%) | 0.7826( 16.13%) | 0.7826( 19.99%) | 0.4565( 19.23%) | - | 0.4222 | ||
TagMyNews | 10% | 0.4500( 0.00%) | 0.4000( 14.29%) | 0.4000( 0.00%) | 0.4000( 11.11%) | 0.4000( 11.11%) | 0.4000( 0.00%) | 0.5000 | |
30% | 0.5556( 16.66%) | 0.5278( 20.83%) | 0.6389( 27.78%) | 0.6389( 0.00%) | 0.6389( 4.17%) | 0.7222( 0.00%) | 0.5833 | ||
50% | 0.8889( 3.23%) | 0.8056( 6.45%) | 0.7778( 0.00%) | 0.9167( 6.46%) | 0.9167( 6.46%) | 0.8611( 3.34%) | 0.7222 | ||
Sentiment Analysis | 10% | 0.4000( 20.01%) | 0.4000( 25.00%) | 0.4000( 20.01%) | 0.4000( 20.01%) | 0.4000( 20.01%) | 0.4667( 74.99%) | 0.3333 | |
30% | 0.8667( 30.00%) | 0.8667( 18.19%) | 0.8000( 9.10%) | 0.8667( 30.00%) | 0.8667( 30.00%) | 0.6667( 100.03%) | 0.5333 | ||
50% | 0.9333( 0.00%) | 0.9333( 6.67%) | 0.9333( 0.00%) | 0.9333( 0.00%) | 0.9333( 0.00%) | 0.7333( 37.50%) | 0.5333 | ||
Average | 0.6619( 6.46%) | 0.6392( 0.79%) | 0.6403( 17.09%) | 0.6687( 7.43%) | 0.5724( 4.52%) | 0.6417( 35.98%) | 0.4555 |

### v2

@@ 3.6. Configurations
Configuration of fault detection methods. NNS and TestRank need to compute the distance between data samples and require the representation of the samples. In this work, we directly use the input embeddings extracted from LLMs as the representation. Besides, TestRank is a learning-based method where training data is required for fault detection. For the clone detection, problem classification, and TagMyNews tasks, we randomly select 200 data from the original datasets as the training data. For the sentiment analysis dataset, as the original dataset only contains 200 data samples and 150 of them are used as test data, we collect the remaining 50 samples as the training data. For the data labeling process, we set the labeling budgets from 10% to 90% with an interval of 10%.
Configuration of LLMs. In our experiments, we set top_p=0.95 for open-sourced LLaMA3 and DeepSeek-Coder, and top_p=1 for closure GPT3.5 and GPT4 which are the default settings used by LLMs. top_p controls the accumulation of token probabilities to select the next word for all LLMs. We set the window size of the token context as max_length= 10k for LLaMA3 and DeepSeek-Coder to match our test data’s token length. The window sizes of GPT3.5 and GPT4 are set as 16385 and 128000 respectively, using the default settings.
@@ 3.7. Implementation and Environment
For the fault detection methods, we reuse the official implementations provided by the original papers and modify them to fit our tasks. For the LLaMA and CodeLLaMA models, we use the resources provided by Hugging Face333Hugging Face, https://huggingface.co/ in this work. For GPT3.5 and GPT4 models, we use the OpenAI’s API to access the models and collect the results. In total, it costs 267 US dollars to run GPT-related experiments. We conduct all LLaMA-related experiments on a high-performance computer cluster. Each cluster node runs a 2.6 GHz Intel Xeon Gold 6132 CPU with an NVIDIA Tesla V100 16G SXM2 GPU.
@@ 4.1. RQ1: Prediction Confidence
First, as most fault detection methods rely on prediction confidence, we check whether LLMs can confidently predict the data in our studied tasks. Figure 2 depicts the distribution (number of data in each confidence interval) of the prediction confidence produced by LLMs, where we set the number of intervals as 30. The first conclusion is that GPT4 has the most confidence in its predictions compared to other LLMs. On average, the confidence of DeepSeekCoder, LLaMA3, GPT3.5, and GPT4 on these four datasets are 81.72%, 59.15%, 77.01%, and 87.27%, respectively. Comparing the accuracy of LLMs and their average confidence (which is a notion of miscalibration (Guo et al., 2017)), we found that our LLMs are not well-calibrated to these four tasks. For example, for DeepSeekCoder on the clone detection task, the accuracy and average confidence are 69.4% and 94.61%, respectively. This means DeepSeekCoder is overconfident with its prediction even though it has poor performance. Overall, except for TagMyNews, LLMs are overconfident in the other three tasks with 16.78% higher confidence than the true accuracy. Additionally, the results show that prediction confidence is concentrated in certain ranges for many cases. For example, for the clone detection - GPT3.5, confidence only falls into four ranges. This means the model has similar confidence to many different inputs, in turn, the output-based fault detection methods (e.g., MaxP) could produce similar uncertainty scores to these inputs.
Interestingly, we found that GPT4 has perfect accuracy (100%) on the problem classification task with high average confidence (0.9827 on average) and a low ECE score of 0.0143. We conjecture there is a data leakage problem of our test data, i.e., our used problem classification dataset is a part of the training data of GPT4. Investigating the exact reason behind this could be future work to better understand what has been learned by the pre-trained LLMs. As there are no faults in GPT4 for the problem classification task, we exclude this model from the following fault detection study.
@@ 4.2. RQ2: Effectiveness of Fault Detection
In this RQ, we explore the performance of existing fault detection methods for LLMs. As mentioned in Section 3.2, we exclude models that have perfect performance (100% accuracy) on the studied tasks.
| Random | Gini | Entropy | Mcp | Maxp | Margin | ATS | NNS | TestRank | |
| DeepSeekCoder-CD | 0.5301 | 0.6586 | 0.6973 | 0.6746 | 0.6586 | 0.6586 | - | 0.6732 | 0.6106 |
| GPT3.5-CD | 0.5121 | 0.6202 | 0.6132 | 0.5774 | 0.6202 | 0.6202 | - | - | - |
| GPT4-CD | 0.5501 | 0.6777 | 0.6535 | 0.6371 | 0.5424 | 0.6777 | - | - | - |
| DeepSeekCoder-PC | 0.6614 | 0.6609 | 0.6796 | 0.7116 | 0.6631 | 0.6670 | 0.7031 | 0.5754 | 0.7169 |
| GPT3.5-PC | 0.7160 | 0.7755 | 0.6984 | 0.7110 | 0.7774 | 0.7811 | 0.7134 | - | - |
| LLaMA3-TMN | 0.6191 | 0.7227 | 0.7569 | 0.6419 | 0.7627 | 0.7737 | 0.7446 | 0.6541 | 0.6683 |
| GPT3.5-TMN | 0.5889 | 0.6620 | 0.6551 | 0.6139 | 0.6903 | 0.7042 | 0.7096 | - | - |
| GPT4-TMN | 0.5395 | 0.8031 | 0.7889 | 0.7265 | 0.7815 | 0.7784 | 0.7944 | - | - |
| LLaMA3-Sentiment | 0.5392 | 0.6490 | 0.6176 | 0.6332 | 0.8131 | 0.9377 | 0.5796 | 0.6147 | 0.5806 |
| GPT3.5-Sentiment | 0.5177 | 0.5852 | 0.5951 | 0.5728 | 0.6305 | 0.6543 | 0.4609 | - | - |
| GPT4-Sentiment | 0.5259 | 0.8000 | 0.8667 | 0.8370 | 0.8000 | 0.8000 | 0.5259 | - | - |
| Average | 0.5727 | 0.6923 | 0.6929 | 0.6670 | 0.7036 | 0.7321 | 0.6540 | 0.6294 | 0.6441 |
Figure 3 depicts the TRC values achieved by fault detection methods in each labeling budget, and Table 2 summarizes the averaged TRC values across all considered labeling budgets. First, considering the averaged results, we can see that all methods have relatively better results than random selection, which indicates the potential of their usefulness. However, we found that the best averaged TRC value on LLMs is only 0.7321, which is far away from the idea detection performance, TRC = 1. Worsely, in some cases, e.g., Sentiment-LLaMA3, some methods have less than 0.2 TRC score, which never happened in conventional deep learning models (LI et al., 2021). This means fault detection for LLMs is a more challenging problem compared to fault detection for conventional deep learning models.
Then, comparing each method, interestingly, we found that the simple method, Margin, which just utilizes the top-1 and top-2 output probabilities to measure the uncertainty, performs the best among our considered methods. This phenomenon is similar to the findings by existing works (Weiss and Tonella, 2022) which stated that simple methods perform the best for neural network test prioritization. Specifically, in more than half of cases (6 out of 11 cases), Margin achieved the highest averaged TRC scores. Methods that require embedding information for the data selection (NNS and TestRank) cannot stand out in LLMs. The reason could be that the embeddings extracted by LLMs are too high-dimensional and diverse to be learned by simple clustering methods (used by NNS) and graph neural networks (used by TestRank).
Answer to RQ2: Under our studied objects, the simple method, Margin, performs the best in fault detection for LLMs. However, there is still a big room for improvement, and more effective methods are needed.
5. MuCS: Mutation-Based Confidence Smoothing
@@
As analyzed in Section 4.1, the concentrated confidence issue can harm the performance of output-based fault detection methods (those methods will give similar uncertainty scores to the inputs), we need to diversify the prediction confidence of LLMs to enhance fault detection. It is challenging since we cannot get the internal information (e.g., output logits) of closed-source LLMs to smooth the prediction confidence. Inspired by previous works (Wang et al., 2021; Li et al., 2023b) that showed the potential of testing deep learning models by mutation analysis (one uses mutation killing score to detect faults and the other one uses mutation analysis to improve the performance of code models), we propose MuCS, a simple yet effective black-box framework to enhance existing fault detection methods by smoothing the prediction confidence from the input-level using prompt mutation.
@@
Figure 4 presents the overall workflow of our proposed framework. It has two main steps, 1) prompt mutation, and 2) confidence smooth. Specifically, given an LLM and an input prompt p, we first utilize transformation methods to mutate the inputs and generate n mutated prompts, , , …, . For NLP tasks, we use text augmentation methods (Marivate and Sefara, 2020) (e.g., token random deletion) to transform prompts into new ones. For the code tasks, we consider both text augmentation methods and code refactoring methods (Al Dallal and Abdin, 2017) (e.g., add new Print() functions) for the prompt mutation. After that, we feed all the mutants to the LLM and collect the output confidence , ,…, produced by the LLM. Finally, a confidence smooth method is employed to compute the final confidence of the LLM on the input. In this work, we use the straightforward method, average, to smooth the confidence. Algorithm 1 summarizes our mutation-based confidence smoothing solution. MuCS first randomly selects multiple (defined by the perturbation size ) mutation operators (line 4) to mutate and then generates (lines 4, 5). Then, MuCS collects all the prediction confidence produced by LLMs using all mutants (line 8). Finally, the averaged confidence is returned as the smoothed confidence (lines 10 and 11).
@@ 5.1. Mutation Operators
Two types of mutation operators have been included in the framework, text augmentation methods and code refactoring methods. In total, five text augmentation methods (including Synonym Replacement, Random Deletion, Random Insertion, Random Swap, and Punctuation Insertion), and four code refactoring methods (including Print Adding, Local Variable Adding, Dead If Adding, and Duplication) have been used in MuCS. Given a prompt with the length of works , each mutation operator is defined as follows.
@@ 6.1. Research Question
To investigate whether our proposed framework MuCS can help fault detection for LLMs, we first analyze the effectiveness of existing fault detection with MuCS as a plug-in. Then, we explore the reason why MuCS enhances fault detection by analyzing its influences on the prediction confidence of LLMs. Our exploration plans to answer the following research questions.
RQ3: How effective is MuCS in enhancing fault detection for LLMs? In this research question, we compare the effectiveness of fault detection methods with and without using MuCS to analyze the usefulness of our proposed framework.
RQ4: How does MuCS affect the final outputs of the LLMs? We want to explore how MuCS influences the outputs of LLMs to figure out why it works on fault detection.
@@ 6.3. RQ3: Effectiveness of MuCS
| Gini-M | Entropy-M | MCP-M | MaxP-M | Margin-M | ATS-M | NNS-M | TestRank-M | BALD | |
| DeepSeekCoder-CD | 0.6886 (4.55) | 0.6890 (1.20) | 0.6592 (2.28) | 0.6886 (4.55) | 0.6886 (4.55) | - | 0.6576 (2.32) | 0.6134 (0.46) | 0.7067 |
| GPT3.5-CD | 0.6873 (10.81) | 0.5989 (2.33) | 0.6369 (10.31) | 0.6873 (10.81) | 0.6873 (10.81) | - | - | - | 0.5827 |
| GPT4-CD | 0.7550 (11.41) | 0.7211 (10.34) | 0.7318 (14.87) | 0.7550 (39.20) | 0.5107 (24.64) | - | - | - | 0.4380 |
| DeepSeekCoder-PC | 0.7462 (12.91) | 0.7134 (4.97) | 0.6883 (3.27) | 0.7412 (11.77) | 0.6516 (2.31) | 0.7212 (2.57) | 0.6984 (21.37) | 0.7425 (3.57) | 0.7145 |
| GPT3.5-PC | 0.8344 (7.59) | 0.8346 (19.49) | 0.7116 (0.08) | 0.8492 (9.24) | 0.8343 (6.81) | 0.7930 (11.16) | - | - | 0.8389 |
| LLaMA3-TMN | 0.7637 (5.68) | 0.7722 (2.02) | 0.7073 (10.18) | 0.7702 (0.99) | 0.7653 (1.09) | 0.7162 (3.82) | 0.7156 (9.41) | 0.7086 (6.03) | 0.7434 |
| GPT3.5-TMN | 0.7431 (12.24) | 0.7319 (11.72) | 0.6208 (1.13) | 0.7630 (10.54) | 0.7481 (6.24) | 0.7708 (8.63) | - | - | 0.7407 |
| GPT4-TMN | 0.7691 (4.23) | 0.7327 (7.12) | 0.7451 (2.55) | 0.7790 (0.32) | 0.7821 (0.48) | 0.7852 (1.16) | - | - | 0.7037 |
| LLaMA3-Sentiment | 0.6417 (1.13) | 0.5899 (4.48) | 0.8423 (33.02) | 0.7587 (6.70) | 0.8643 (7.83) | 0.6211 (7.16) | 0.6006 (2.30) | 0.6065 (4.46) | 0.6199 |
| GPT3.5-Sentiment | 0.7654 (30.80) | 0.7646 (28.49) | 0.7959 (38.94) | 0.7539 (19.58) | 0.6782 (3.65) | 0.7860 (70.53) | - | - | 0.6379 |
| GPT4-Sentiment | 0.8593 (7.41) | 0.8593 (0.85) | 0.8444 (0.88) | 0.8593 (7.41) | 0.8519 (6.49) | 0.7481 (42.24) | - | - | 0.6667 |
| Average | 0.7503 (8.39) | 0.7280 (5.05) | 0.7258 (8.81) | 0.7641 (8.60) | 0.7329 (0.12) | 0.7427 (13.57) | 0.6681 (6.15) | 0.6678 (3.67) | 0.6720 |
| Gini | Entropy | Mcp | Maxp | Margin | ATS | NNS | TestRank | ||
| Welch’s t-test | Significance | 2.8873 | 1.7373 | 3.0355 | 2.5787 | 1.1567 | 3.1049 | 1.1405 | 0.7018 |
| P-value | 4.10E-03 | 8.31E-02 | 2.56E-03 | 1.03E-02 | 2.48E-01 | 2.11E-03 | 2.56E-01 | 4.84E-01 | |
| Wilcoxon signed-rank test | Significance | 500.5 | 1136.5 | 770.5 | 663.5 | 1273.5 | 283 | 148.5 | 42.5 |
| P-value | 4.82E-10 | 2.47E-04 | 1.30E-06 | 9.15E-09 | 1.18E-03 | 1.95E-06 | 1.09E-02 | 2.11E-05 |